OpenStudy (taylovesmatthew):

simplify the expression. (2n^7)4? i don't understand this at all to even try and figure it out

OpenStudy (anonymous):

\[\large (2n^4)^7=2^4n^{4\times 7}\] is a start

OpenStudy (taylovesmatthew):

@satellite73 i don't know what to do still..